# The radiation condition for Helmholtz equations above locally perturbed periodic surfaces

### v1

@@ 6.1 Decay of $u_{0}$
Since we will need the decay rate of $u_{0}$ in the strong form later, with the interior regularity property for elliptic equations (see [14]), we have the following corollary.
@@ 6.3 Decay of $u_{2}$
For term (I), from integration by parts,
$(I)$ $=\int_{0}^{\delta}sw(s+\alpha_{0},x_{0})e^{\mathrm{i}2\pi js}\,\mathrm{d}s\,=\frac{1}{\mathrm{i}2\pi j}\int_{0}^{\delta}s\phi(s+\alpha_{0})\,\mathrm{d}e\,^{\mathrm{i}2\pi js}$
$=-\frac{1}{\mathrm{i}2\pi j}\int_{0}^{\delta}(w(s+\alpha_{0},x_{0})+sw^{\prime}(s+\alpha_{0},x_{0}))e^{\mathrm{i}2\pi js}\,\mathrm{d}s\,$
$=-\frac{1}{4\pi^{2}j^{2}}\int_{0}^{\delta}(w(s+\alpha_{0},x_{0})+sw^{\prime}(s+\alpha_{0},x_{0}))\,\mathrm{d}e\,^{\mathrm{i}2\pi js}$
$=\frac{1}{4\pi^{2}j^{2}}w(\alpha_{0},x_{0})+\frac{1}{4\pi^{2}j^{2}}\int_{0}^{\delta}(w(s+\alpha_{0},x_{0})+sw^{\prime}(s+\alpha_{0},x_{0}))^{\prime}e^{\mathrm{i}2\pi js}\,\mathrm{d}\xi\,.$
Thus there is a constant $C>0$ such that
@@ 7 Radiation condition for $v_{0}$
In this section, we will prove Part (B) in Theorem 4.2 for $v_{0}$, which is defined by (4.7). We consider the radiation condition for the following layer potential (for simplicity, denote $u=v_{0}$ and $\phi=u_{0}\big{|}_{\Gamma^{H}}$ in the definition):
@@
where $H_{0}^{(1)}(t)$ is the Hankel function of the first kind. With (11.11),
@@
Then with (11.12), and for simplicity let $\widetilde{x}:=x-(0,H)^{\top}$,
@@
We first use the asymptotic behaviour (11.14) for $n=0,1$ and $\ell=1$:
@@
In order to obtain a more explicit analysis, we need to consider the term
@@
particularly. This is concluded in the following lemma.
@@
$|\widetilde{x}-\widetilde{y}|\geq|\widetilde{x}|-|\widetilde{x}|\geq|\widetilde{x}|-2\sqrt{|\widetilde{x}|}\geq\frac{r}{2}.$
@@
$|\widetilde{x}||\widetilde{x}-\widetilde{y}|=|\widetilde{x}|^{2}\sqrt{1+t}\leq|\widetilde{x}|^{2}\left(1+\frac{1}{2}t\right)=|\widetilde{x}|^{2}+\frac{|y|^{2}}{2}-\left<x,y\right>,$
@@
Similar as the proof of Theorem 7.2,
@@ 8 Radiation condition for $v_{1}$
  • In Section 8.2, we consider Case II: $|\alpha_{0}+j|>k$. These $j$’s define a set $J_{+}$.
  • In Section 8.3, we consider Case III: $|\alpha_{0}+j|=k$. These $j$’s define a set $J_{0}$.
@@
  • In Case I, $|\alpha+j|<k-\delta$ for all $\alpha\in[\alpha_{0}-\delta,\alpha_{0}+\delta]$.
  • In Case II, $|\alpha+j|>k+\delta$ for all $\alpha\in[\alpha_{0}-\delta,\alpha_{0}+\delta]$.
@@
As described in the beginning of this section, for any $\alpha_{0}\in S$, $|\alpha_{0}+j|<k$ for all $j\in J_{-}(\alpha_{0})$, which is a finite set. Define
@@
Given any $\psi\in C_{0}^{\infty}(\beta_{1},\beta_{2})$, there is a constant $C>0$ which only depens on $\psi$ such that
@@
Since $\theta-\theta_{*}\in(\delta_{0}-\pi,\pi-\delta_{0})$, the map $\theta\mapsto\sin\left(\frac{\theta-\theta_{*}}{2}\right)$ is an injection and anaytic. Let $t:=\sin\left(\frac{\theta-\theta_{*}}{2}\right)\in[\gamma_{1},\gamma_{2}]$, then
@@
$\zeta(t)=\psi\left(2\arcsin t+\theta_{*}\right)\frac{2}{\sqrt{1-t^{2}}}\sqrt{\frac{2\arcsin t+\theta_{*}-\theta_{0}}{t-a}}\in C_{0}^{\infty}(\gamma_{1},\gamma_{2}).$
Thus $v$ has exactly the same form as the integral $I_{2}^{C}(r)$ defined by (5.12). From Corollary (5.17), there is a constant $C>0$ such that
@@
which has exactly the same as $I_{1}(r)$ defined in (5.1). From Theorem 5.1, $|I_{1}(r)|\leq Cr^{-1/2}$. From the previous computation, we get that
@@
$\zeta(t)=-\mathrm{i}k^{5/2}\frac{4}{\sqrt{1-t^{2}}}\sqrt{\frac{2\arcsin t+\theta_{*}-\theta_{0}}{t-a}}\psi\left(\arcsin(2t)+\theta_{*}\right)\in C_{0}^{\infty}(\gamma_{1},\gamma_{2})$
Although it has exactly the same form as $I_{2}^{C}(r)$ as defined in (5.12), we still need a further study since the decay rate is too low compared to our expectation. With integration by parts,
@@
$=\frac{1}{4\mathrm{i}kr}\int_{\gamma_{1}}^{\gamma_{2}}\left[\frac{t}{2\sqrt{t-a}}\zeta(t)+\sqrt{t-a}\zeta(t)+t\sqrt{t-a}\zeta^{\prime}(t)\right]e^{-2\mathrm{i}krt^{2}}\,\mathrm{d}t\,$
@@
Oberserve that (i) has exactly the form of $I_{1}^{C}(r)$ defined in (5.11) and (ii) has the form of $I_{2}^{C}(r)$ defined in (5.12). From Theorem 5.7 and Corollary 5.8, there is a constant $C>0$ such that
@@
The estimation holds similarly for $\frac{\partial v}{\partial r}-\mathrm{i}kv$. The proof is finished. ∎
@@
Since $x_{2}-H\leq\frac{r}{2}$, then $|x_{1}|\geq\frac{\sqrt{3}}{2}r>\frac{r}{2}$. Note that both (8.7) and (8.8) have exactly the same form as $I_{2}^{A}(r)$ defined by (5.3), then from Corollary 5.3,
@@
where $C$ is determined by $\left\|\phi(\alpha)e^{-\sqrt{(\alpha+j)^{2}-k^{2}}(x_{2}-H)}\right\|_{C^{3}(\mathbb{R})}$.
The key step in the estimation lies in the fact that $\left\|\phi(\alpha)e^{-\sqrt{(\alpha+j)^{2}-k^{2}}(x_{2}-H)}\right\|_{C^{3}(\mathbb{R})}$ also depends on $x_{2}$, which is not bounded. Note that the leading term with respect to $x_{2}-H$ has the form of
@@
In this section we consider the particular case that $|\alpha_{0}+j|=k$ where $j\in\mathbb{Z}$. For $\alpha\in(\alpha_{0}-\delta,\alpha_{0}+\delta)$, either $|\alpha+j|\leq k$ or $|\alpha+j|\geq k$. Since both cases are complex and the proofs are long, we will discuss in two subsections. For simplicity, we assume that $\alpha_{0}+j=k$, then (i) when $\alpha\in(\alpha_{0}-\delta,\alpha_{0}]$, $\alpha+j\leq k$ (see Section 8.3.1), (ii) when $\alpha\in[\alpha_{0},\alpha_{0}+\delta)$, $\alpha+j\geq k$ (see Section 8.3.2). The case $\alpha_{0}+j=-k$ is very similar thus we will not discuss in this paper.
@@ 8.3.1 Case (i): $\alpha\in(\alpha_{0}-\delta,\alpha_{0}]$
Similar to (8.1) (with a little difference in the definition of $\phi$), with the polar coordinate,
@@
$\psi(\theta):=\mathrm{i}\sqrt{2}\phi(k\cos\theta-j)\frac{\sin(\theta/2)}{\sin\theta}\in C^{\infty}[0,\beta)$
@@
i) First consider the case that $\theta_{*}\in(2\beta,\pi-2\beta)$, which implies that $\sin(\theta-\theta_{*})\geq\sin(\beta)>0$ holds uniformly for $\theta\in[0,\beta]$. In this case, with integration by parts,
@@
ii) Second consider the case that $\theta_{*}\in[0,2\beta]$. The technique of integration by parts now fails since $\sin(\theta-\theta_{*})\rightarrow 0$ when $\theta\rightarrow\theta_{*}$. In this case, let
@@
$\sin\left(\frac{\theta-\theta_{*}}{2}\right)+\sin\left(\frac{\theta_{*}}{2}\right)=2\sin\left(\frac{\theta}{4}\right)\cos\left(\frac{\theta-2\theta_{*}}{4}\right),$
@@
$\frac{\sin\left(\theta\right)}{\sin\left(\frac{\theta-\theta_{*}}{2}\right)+\sin\left(\frac{\theta_{*}}{2}\right)}=\frac{\sin\left(\theta\right)}{\sin\left(\theta/4\right)}\frac{1}{2\cos((\theta-2\theta_{*})/4)}\in C^{\infty}[0,2\beta].$
@@
Note that now $v$ has exactly the same form of $I_{B}^{2}(r)$ defined by (5.9). Using Theorem 5.6, there is a constant $C>0$ such that
@@
Again the form is similar to $I_{B}^{2}(r)$ thus we can use Theorem 5.6. Thus we have a constant $C>0$ such that
@@
i) First consider $\theta_{*}\in(2\beta,\pi-2\beta)$, where the decay rate $O(r^{-2})$ comes directly from integration by parts, following the proof of Theorem 8.6.
@@
$\zeta_{2}(\theta)$ $=-(\cos(\theta+\tau_{*})+1)\psi(\theta)$
@@
$(I_{2})$ $=\frac{e^{-\mathrm{i}kr}}{4\mathrm{i}kr}\int_{\gamma_{1}}^{\gamma_{1}}\frac{1}{t}\lambda(t)\rho_{1}(t)\,\mathrm{d}e\,^{2\mathrm{i}krt^{2}}$
@@
Since $\gamma_{1}\geq r^{-1/2}$ and $\gamma_{2}>\gamma_{1}$, we can estimate immediately that
@@
It means that when $\tau_{*}$ is close to 00, the value will be very large thus $\lambda^{\prime}$ is not uniformly bounded. But note that $\eta^{\prime}(\theta)>0$ for any $\theta\geq 0$, it implies that $\lambda$ is strictly increasing. Thus
@@ 8.3.2 Case (ii): $\alpha\in[\alpha_{0},\alpha_{0}+\delta)$
where $\psi_{1}$ and $\psi_{2}$ only depend on $\theta_{*}$ and $k$. Both functions are $C^{\infty}[0,\delta)$ and vanishes around $\delta$. Note that $v$ defined by (8.15) has exactly the same form as $(I)$ so we will not discuss this case separately.
@@
$|(Iii)|\leq C\frac{1}{x_{1}}\quad\Rightarrow\quad|(I)|\leq Cr^{-1/2}$
@@
which vanish around $\delta$. Now we consdier $(II_{1})$. Again with integration by parts and the inequality $\sqrt{k^{2}+2kt}\geq\sqrt{2kt}$,
$|(II_{1})|$ $=\left|\frac{1}{\mathrm{i}x_{1}}\int_{0}^{\delta}\zeta_{3}(t)e^{-\xi\sqrt{t^{2}+2kt}}\,\mathrm{d}e\,^{\mathrm{i}tx_{1}}\right|$
$\leq\left|\frac{1}{\mathrm{i}x_{1}}\zeta_{3}(0)\right|+\left|\frac{\mathrm{i}}{x_{1}}\int_{0}^{\delta}\left[\zeta^{\prime}_{3}(t)-\frac{\xi(t+k)}{\sqrt{t^{2}+2kt}}\right]e^{\mathrm{i}tx_{1}-\xi\sqrt{k^{2}+2kt}}\,\mathrm{d}t\,\right|$
$\leq Cr^{-1}+Cr^{-1}+Cr^{-1}\int_{0}^{\delta}\frac{\xi}{\sqrt{2kt}}e^{-\xi\sqrt{2kt}}\,\mathrm{d}t\,\leq Cr^{-1}.$
For $(II_{2})$, with similar technique,

### v2

@@ 5.1 Decay of $u_{0}$
Since we will need the decay rate of $u_{0}\big{|}_{\Gamma^{H}}$ in the strong form later, with the interior regularity property for elliptic equations (see [13]), we have the following corollary.
@@ 5.3 Decay of $u_{2}$
From Theorem B.4, there is a constant $C>0$ such that
@@ 6 Radiation condition for $v_{0}$
In this section, we will prove Part (B) in Theorem 4.2 for $v_{0}$, which is defined by (4.6). We consider the radiation condition for the following layer potential (for simplicity, denote $u=v_{0}$ and $\phi=2u_{0}\big{|}_{\Gamma^{H}}$ in the definition):
@@
Remark 6.1.
The proof mainly follows the proofs in [17, 19]. The only difference is the decay rate of the density function $\phi$, which is super-algebraic in this paper, but the density function in these references decay at the rate of $(1+|y|^{2})^{-3/4}$.
@@
where $H_{0}^{(1)}(t)$ is the Hankel function of the first kind. With (A.5),
@@
Then with (A.6), and for simplicity let $\widetilde{x}:=x-(0,H)^{\top}$,
@@
We first use the asymptotic behaviour (A.7) for $n=0,1$ and $\ell=1$:
@@
Now we will focus on the term
@@
particularly. The property of this term is concluded in the following lemma.
@@
$|\widetilde{x}-\widetilde{y}|\geq|\widetilde{x}|-|\widetilde{y}|>|\widetilde{x}|-2\sqrt{|\widetilde{x}|}\geq\frac{r}{2}.$
@@
$|\widetilde{x}||\widetilde{x}-\widetilde{y}|=|\widetilde{x}|^{2}\sqrt{1+t}\leq|\widetilde{x}|^{2}\left(1+\frac{1}{2}t\right)=|\widetilde{x}|^{2}+\frac{|\widetilde{y}|^{2}}{2}-\left<\widetilde{x},\widetilde{y}\right>,$
@@
Similar as the proof of Theorem 6.3, we split the integral into two parts:
@@ 7 Radiation condition for $v_{1}$
  • In Section 7.1, we consider Case I: $j\in J_{-}$, i.e., $|\alpha_{0}+j|<k$.
  • In Section 7.2, we consider Case II: $j\in J_{+}$, i.e., $|\alpha_{0}+j|>k$.
  • In Section 7.3, we consider Case III: $j\in J_{0}$, i.e., $|\alpha_{0}+j|=k$.
@@
  • In Case I, $|\alpha+j|<k-\delta$ for all $\alpha\in[\alpha_{0}-\delta,\alpha_{0}+\delta]\subset\Lambda$.
  • In Case II, $|\alpha+j|>k+\delta$ for all $\alpha\in[\alpha_{0}-\delta,\alpha_{0}+\delta]\subset\Lambda$.
@@
As described in the beginning of this section, for any fixed $\alpha_{0}\in S$, $|\alpha_{0}+j|<k$ for all $j\in J_{-}(\alpha_{0})$, which is a finite set. Define
@@
Given any $\psi\in C_{0}^{\infty}(\beta_{1},\beta_{2})$ and let $v$ be defined by (7.2). There is a constant $C>0$ which only depens on $\psi$ such that
@@
Since $\theta-\theta_{*}\in(\delta_{0}-\pi,\pi-\delta_{0})$, the map $\theta\mapsto\sin\left(\frac{\theta-\theta_{*}}{2}\right)$ is an injection and anaytic. Let $t:=\sin\left(\frac{\theta-\theta_{*}}{2}\right)\in[\gamma_{1},\gamma_{2}]\subset(-1,1)$, then
@@
$\zeta(t)=\psi\left(2\arcsin t+\theta_{*}\right)\frac{2}{\sqrt{1-t^{2}}}\sqrt{\frac{2\arcsin t-2\arcsin a}{t-a}}\in C_{0}^{\infty}(\gamma_{1},\gamma_{2}).$
Thus $v$ has exactly the same form as the integral $I_{2}^{C}(r)$ defined by (B.12). From Corollary (B.17), there is a constant $C>0$ such that
@@
which has exactly the same as $I_{1}(r)$ defined in (B.1). From Theorem B.1, $|I_{1}(r)|\leq Cr^{-1/2}$. From the previous computation, we get that
@@
$\zeta(t)=-\mathrm{i}k^{5/2}\frac{4}{\sqrt{1-t^{2}}}\sqrt{\frac{2\arcsin t-2\arcsin a}{t-a}}\psi\left(\arcsin(2t)+\theta_{*}\right)\in C_{0}^{\infty}(\gamma_{1},\gamma_{2})$
Although it has exactly the same form as $I_{2}^{C}(r)$ as defined in (B.12), we still need a further study since the decay rate is too low compared to our expectation. With integration by parts,
@@
Oberserve that (i) has exactly the form of $I_{1}^{C}(r)$ defined in (B.11) and (ii) has the form of $I_{2}^{C}(r)$ defined in (B.12). From Theorem B.7 and Corollary B.8, there is a constant $C>0$ such that
@@
The estimation holds similarly for $\frac{\partial v}{\partial r}-\mathrm{i}kv$. The only difference is that $|\psi|\sim|j|$, which can be merged into the exponential term. The proof is finished. ∎
@@
Since $x_{2}-H\leq\frac{r}{2}$, then $|x_{1}|\geq\frac{\sqrt{3}}{2}r>\frac{r}{2}$. Note that both (7.7) and (7.8) have exactly the same form as $I_{2}^{A}(r)$ defined by (B.3), then from Corollary B.3,
@@
where $C$ is determined by $\left\|\phi(\alpha)e^{-\sqrt{(\alpha+j)^{2}-k^{2}}(x_{2}-H)}\right\|_{C^{3}(\mathbb{R})}$, where the norm of $C^{3}(\mathbb{R})$ is computed with respect to $\alpha$ with fixed $x$.
Note that the fact that $\left\|\phi(\alpha)e^{-\sqrt{(\alpha+j)^{2}-k^{2}}(x_{2}-H)}\right\|_{C^{3}(\mathbb{R})}$ also depends on $x_{2}$, which is not bounded when $r\rightarrow\infty$. Note that the leading term with respect to $x_{2}-H$ has the form of
@@
In this section we consider the particular case that $|\alpha_{0}+j|=k$ where $j\in\mathbb{Z}$. For $\alpha\in(\alpha_{0}-\delta,\alpha_{0}+\delta)$, either $|\alpha+j|\leq k$ or $|\alpha+j|\geq k$. Since both cases are complex and the proofs are long, we will discuss them in two subsections. For simplicity, we assume that $\alpha_{0}+j=k$, then (i) when $\alpha\in(\alpha_{0}-\delta,\alpha_{0}]$, $\alpha+j\leq k$ (see Section 7.3.1), (ii) when $\alpha\in[\alpha_{0},\alpha_{0}+\delta)$, $\alpha+j\geq k$ (see Section 7.3.2). The case $\alpha_{0}+j=-k$ is very similar thus we will not discuss in this paper.
@@ 7.3.1 Case (i): $\alpha\in(\alpha_{0}-\delta,\alpha_{0}]$
Similar to (7.1), with the polar coordinate and a slightly different definition of $\phi$,
@@
$\psi(\theta):=\mathrm{i}\sqrt{2}\phi(k\cos\theta-j)\frac{\sin^{2}(\theta/2)}{\sin^{2}\theta}\in C^{\infty}[0,\beta)$
@@
i) First consider the case that $\theta_{*}\in(2\beta,\pi-2\beta)$, which implies that $\left|\sin(\theta-\theta_{*})\right|\geq\sin(\beta)>0$ holds uniformly for $\theta\in[0,\beta]$. In this case, with integration by parts,
@@
ii) Second consider the case that $\theta_{*}\in[0,2\beta]$. The technique of integration by parts now fails since $\sin(\theta-\theta_{*})$ is close/equal to zero since $|\theta-\theta_{*}|$ can be very small. In this case, let
@@
$t+a=\sin\left(\frac{\theta-\theta_{*}}{2}\right)+\sin\left(\frac{\theta_{*}}{2}\right)=2\sin\left(\frac{\theta}{4}\right)\cos\left(\frac{\theta-2\theta_{*}}{4}\right),$
@@
$\frac{\sin(\theta)}{t+a}=\frac{\sin\left(\theta\right)}{\sin\left(\frac{\theta-\theta_{*}}{2}\right)+\sin\left(\frac{\theta_{*}}{2}\right)}=\frac{\sin\left(\theta\right)}{\sin\left(\theta/4\right)}\frac{1}{2\cos((\theta-2\theta_{*})/4)}\in C^{\infty}[0,2\beta].$
@@
Note that now $v$ has exactly the same form of $I_{B}^{2}(r)$ defined by (B.9). Using Theorem B.6, there is a constant $C>0$ such that
@@
Again the form is the same as $I_{B}^{2}(r)$ thus we can use Theorem B.6. Thus we have a constant $C>0$ such that
@@
i) First consider $\theta_{*}\in(2\beta,\pi-2\beta)$, where the decay rate $O(r^{-2})$ comes directly from integration by parts, following the proof of i) Theorem 7.6.
@@
$\zeta_{2}(\theta)$ $=-\frac{\sin^{2}(\theta+\tau_{*})}{2\sin^{2}((\theta+\tau_{*})/2)}\psi(\theta)$
@@
$(I_{2})$ $=\frac{e^{-\mathrm{i}kr}}{4\mathrm{i}kr}\int_{\gamma_{1}}^{\gamma_{2}}\frac{1}{t}\lambda(t)\rho_{1}(t)\,\mathrm{d}e\,^{2\mathrm{i}krt^{2}}$
@@
Since $\gamma_{1}\geq\frac{1}{2}r^{-1/2}$ and $\gamma_{2}>\gamma_{1}$, we can estimate immediately that
@@
It means that when $\tau_{*}$ is close to 00, the value will become very large thus $\lambda^{\prime}$ is not uniformly bounded. But note that $\eta^{\prime}(\theta)>0$ for any $\theta\geq 0$, it implies that $\lambda$ is strictly increasing. Thus
@@ 7.3.2 Case (ii): $\alpha\in[\alpha_{0},\alpha_{0}+\delta)$
where $\psi_{1}$ and $\psi_{2}$ only depend on $\theta_{*}$ and $k$. Both functions are $C^{\infty}[0,\delta)$ and vanishes around $\delta$. Note that $v$ defined by (7.15) has exactly the same form as $(I)$, thus the discussion is included in the proof of $\frac{\partial v}{\partial r}-\mathrm{i}kv$.
@@
$|(I_{2})|\leq C\frac{1}{|x_{1}|}\quad\Rightarrow\quad|(I)|\leq Cr^{-3/2}$
@@
which vanish around $\delta$. For the term $(II_{1})$, since it has exactly the same form of $(I_{2})$, we can easily estimate that
$|(II_{1})|\leq Cr^{-1}.$
For $(II_{2})$, with integration by parts and the inequality $\sqrt{k^{2}+2kt}\geq\sqrt{2kt}$,